OpenStudy (anonymous):

Simplify this equation:

OpenStudy (anonymous):

\[\frac{ d }{ 29f^2 } * \frac{ 2f^3 }{ d^4 }\]

geerky42 (geerky42):

Remember, \(\Large\dfrac{a^m}{a^n} = a^{m-n} = \dfrac{1}{a^{n-m}}\)

OpenStudy (anonymous):

okok, i gotcha. thanks!